OpenStudy (anonymous):

a concession stand vendor counts the money in the register. after 2 h it has $82.75. after 6 h is has $ 360.75. how much money will be in the register at the end of an 8h shift?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

well we can assume it is linear I geuss? Use the points (2,82.75) and (6,360.75) and find the equation of the line through them. Then plug 8 into your equation to estimate the amount at the end of 8 hours, assuming that this relationship is linear

OpenStudy (anonymous):

what would the solution be?

OpenStudy (anonymous):

do you know how to find slope, and use point slope form for the equation of a line

OpenStudy (anonymous):

?

OpenStudy (anonymous):

or just find money per hour: (360.75-82.75)/4=69.5 per hour

OpenStudy (anonymous):

did you get 489.25?

OpenStudy (anonymous):

then in two more hours there will be 139 more. 360.25+139=499.25

OpenStudy (anonymous):

i get 499.25

OpenStudy (anonymous):

woops i made a mistake

OpenStudy (anonymous):

its 360.75+139=499.75

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

no rsvitale, you are incorrect. because initially the vendor had to start at 0,0. and your line doesn't pass through that

OpenStudy (anonymous):

you would start with some money in the register for change

OpenStudy (anonymous):

there was nothing saying you start with no money

OpenStudy (anonymous):

there is nothing to suggest its linear either.

OpenStudy (anonymous):

so there is not really enough information,

OpenStudy (anonymous):

yup, thats my sense too.

OpenStudy (anonymous):

there are a lot of functions that go through 2 points

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

the initial conditions should have been specified, or at least the peak purchase times.

OpenStudy (anonymous):

i assumed since it is algebra 2 it was linear, i don't remember doing anything more advanced than that when I took that class

OpenStudy (anonymous):

oh I see. then it should be all right.

OpenStudy (anonymous):

actually, rsvitale, if you plot the graph, at hour 0, the vendor has negative money. so your argument that he has some money for change is not valid.

OpenStudy (anonymous):

well rsvitale was right because my quiz showed i got that question right

OpenStudy (anonymous):

I am sure he was right. but the problem doesn't give enough information. I think for the class you are taking, you assume that it is linear, but you cannot make that assumption as you study more advanced topics.

OpenStudy (anonymous):

true